A car moving at 43 km/h skids 15 m with locked brakes. How far will the car skid with locked brakes at 86 km/h ?
1
Expert's answer
2018-10-29T16:46:09-0400

s=(v^2-v_0^2)/2a

43 km/h = 11.94 m/sec

a=(v^2-v_0^2)/2s=(0-〖11.94〗^2)/(2·15)=-4.75 meter/〖sec〗^2

If a car moving at 86 km/h

86 km/h=23.89 m/sec

s=(0-〖23.89〗^2)/(2·(-4.75) )=60 metres